# 2cm2cm1cm1cm

### v4

@@ 3.6 Shale-Stinespring-like Condition for Generators of Bogoliubov Transformations
This series is of course well-defined for any $\varphi\in\mathcal{H}$ since $\tilde{V}_{t,s}\varphi$ becomes an absolutely summable series in the Hilbert space $\mathcal{H}$. In fact, $(\tilde{V}_{t,s})_{s,t\in\mathbb{R}}$ is another strongly continuous two-parameter operator family solving the non-autonomous evolution equations (134) and is therefore nothing else that $(V_{t,s})_{s,t\in\mathbb{R}}$, by uniqueness [^25] Compute that, for any $s,t\in\mathbb{R}$ and $\varphi\in\mathcal{H}$, $(\tilde{V}_{t,s}-V_{t,s})\varphi=\int_{s}^{t}\partial_{\tau}\{V_{t,\tau}\tilde{V}_{\tau,s}\}\varphi\ \mathrm{d}\tau=0$. of the solution to (134). This is completly standard. Using the representation (135) of the operator family $(d_{t})_{t\in\mathbb{R}}$ and
@@
Take any self-adjoint operator $h$ satisfying Assumption 1 and use Lemmata 3.13 and 3.14 (i) to represent the group $(\mathrm{e}^{ith})_{t\in\mathbb{R}}$ of Bogoliubov transformations. If $D_{0}\in\mathcal{L}^{2}(\mathfrak{h})$ then $\mathbf{v}_{t}\in\mathcal{L}^{2}(\mathfrak{h})$ for all $t\in\mathbb{R}$.
@@
Take any self-adjoint operator $h$ acting on $\mathcal{H}=\mathfrak{h}\oplus\mathfrak{h}$ satisfying Assumption 1 and use Lemmata 3.13 and 3.14 (i) to represent the group $(\mathrm{e}^{ith})_{t\in\mathbb{R}}$ of Bogoliubov transformations. If $D_{0}\in\mathcal{B}(\mathfrak{h})$ then, for any $t\in\mathbb{R}$,
@@
Proof. Using Lemmata 3.15 and 3.16 with $D_{0}\in\mathcal{B}(\mathfrak{h})$ and $c=\infty$ as well as the mapping (150), we obtain that
@@
Proof. This is an obvious consequence of Lemma 3.21. In particular, since $(\mathbf{v}_{t})_{t\in\left[-c,c\right]}\subseteq\mathcal{L}^{2}(\mathfrak{h})$ is continuous in $\mathcal{L}^{2}(\mathfrak{h})$ at $t=0$ with $\mathbf{v}_{0}=0$, there is $\delta\in\mathbb{R}^{+}$ such that
$\sup_{t\in\left[-\delta,\delta\right]}\left\|\mathbf{v}_{t}\right\|_{2}<+\infty\ .$
Moreover, for the same reason, given any $\varepsilon\in\mathbb{R}^{+}$ there exists $\tilde{\delta}\in\left[-\delta,\delta\right]$ such that
$\sup_{t\in[-\tilde{\delta},\tilde{\delta}]}\left\|\mathbf{v}_{t}\right\|_{2}\leq\frac{\varepsilon}{1+0.5\left\|D_{0}\right\|_{\mathrm{op}}^{2}\delta^{2}}<+\infty\ ,$
which in turns implies that $\left\|\varkappa_{t}(D_{0})\right\|_{2}\leq\varepsilon$ for any $t\in[-\tilde{\delta},\tilde{\delta}]$.
@@
in the Hilbert-Schmidt topology, as $t\rightarrow 0$. For instance, $\mathcal{O}(1)=0$ when $D_{0}\mathrm{e}^{-i\tau\Upsilon_{0}^{\top}/2}=\mathrm{e}^{i\tau\Upsilon_{0}/2}D_{0}$ for any $\tau\in[0,t]$. In the following lemma, we give the simple case for which the operator $D_{0}$ intertwines [^26] We say that an operator $A$ intertwines between two operators $B$ and $C$ when $AB=CA$. between $\Upsilon_{0}$ and $\Upsilon_{0}^{\top}$ (resp. $-\Upsilon_{0}^{\top}$).
@@
Take any self-adjoint operator $h$ acting on $\mathcal{H}=\mathfrak{h}\oplus\mathfrak{h}$ satisfying Assumption 1 with $D_{0}\in\mathcal{B}(\mathfrak{h})$ and use Lemmata 3.13 and 3.14 (i) to represent the group $(\mathrm{e}^{ith})_{t\in\mathbb{R}}$ of Bogoliubov transformations.
@@
    Assume $D_{0}\Upsilon_{0}=-\Upsilon_{0}^{\top}D_{0}$ and $\mathbf{v}_{t}\in\mathcal{L}^{2}(\mathfrak{h})$ for positive times $t$ in a neighborhood of 00. Then, $D_{0}\in\mathcal{L}^{2}(\mathfrak{h})$ and $\mathbf{v}_{t}\in\mathcal{L}^{2}(\mathfrak{h})$ for all $t\in\mathbb{R}$.
@@
    Assume $D_{0}\Upsilon_{0}=\Upsilon_{0}^{\top}D_{0}$ and $\mathbf{v}_{t}\in\mathcal{L}^{2}(\mathfrak{h})$ for positive times $t$ in a neighborhood of 00. If additionally
@@
Step 1: Assume all conditions of the lemma as well as $\Upsilon_{0}D_{0}=\pm D_{0}\Upsilon_{0}^{\top}$. We show here that $D_{0}$ also intertwines between the groups generated by $i\Upsilon_{0}$ and ±$\pm i\Upsilon_{0}^{\top}$. Since $D_{0}\in\mathcal{B}(\mathfrak{h})$, the equality $\Upsilon_{0}D_{0}=\pm D_{0}\Upsilon_{0}^{\top}$ means in particular that
@@
for small times $t\in\left[0,\delta\right]$. In particular, if $D_{0}\Upsilon_{0}=-\Upsilon_{0}^{\top}D_{0}$ and $\mathbf{v}_{t}\in\mathcal{L}^{2}(\mathfrak{h})$ for positive times $t$ in a neighborhood of 00, then $D_{0}\in\mathcal{L}^{2}(\mathfrak{h})$, which, combined with Lemma 3.20, leads to Assertion (i).
@@
In particular, if (155) holds true as well as $D_{0}\Upsilon_{0}=\Upsilon_{0}^{\top}D_{0}$ and $\mathbf{v}_{t}\in\mathcal{L}^{2}(\mathfrak{h})$ for positive times $t$ in a neighborhood of 00, then $D_{0}\in\mathcal{L}^{2}(\mathfrak{h})$, which, combined with Lemma 3.20, leads to Assertion (ii).
@@
$\Upsilon_{0}\in\mathcal{B}(\mathfrak{h})$ implies the norm continuity of the group $(\mathrm{e}^{it\Upsilon_{0}})_{t\in\mathbb{R}}$ of unitary operators, see, e.g., [31, Chapter I, Proposition 3.5 and Theorem 3.7]. Therefore, Lemma 3.24 (ii) clearly holds also true when $\Upsilon_{0}\in\mathcal{B}(\mathfrak{h})$, $D_{0}\Upsilon_{0}=\Upsilon_{0}^{\top}D_{0}$ and $\mathbf{v}_{t}\in\mathcal{L}^{2}(\mathfrak{h})$ for positive times $t$ in a neighborhood of 00.
@@
Take any self-adjoint operator $h$ satisfying Assumption 1 and use Lemmata 3.13 and 3.14 (i) to represent the group $(\mathrm{e}^{ith})_{t\in\mathbb{R}}$ of Bogoliubov transformations. Assume that $(\mathrm{e}^{ith})_{t\in\mathbb{R}}$ can be implemented in the Fock space $\mathcal{F}$ via a strongly continuous group $(\mathrm{e}^{it\mathrm{H}/2})_{t\in\mathbb{R}}$, $\mathrm{H}=\mathrm{H}^{\ast}$ being a self-adjoint operator acting on $\mathcal{F}$, as in Definition 2.10. Then the following assertions hold true:
@@
The Fock space representation $\mathbf{\pi}_{\mathcal{F}}$ is irreducible, see, e.g. [32, Definition 2.3.7] and [27, Proposition 5.2.2]. It follows from Schur’s lemma [52, Proposition 9.20] that
[…]
@@
($-\partial_{t}^{2}f\left(0\right)$ is in fact the quantum fluctuation of the energy observable $\mathrm{H}$ in the state $\left\langle\Psi,\left(\cdot\right)\Psi\right\rangle_{\mathfrak{h}}$, which is positive by the Cauchy-Schwarz inequality.) As $f\left(t\right)>0$ at least for all $t\in\left[-\delta,\delta\right]$, it follows from (178) that, for any $t\in\left[-\delta,\delta\right]$,
@@
Proof. Combine Proposition 3.26 with Lemma 3.18 as well as Corollaries 3.22 and 3.23.
3.7 Paradigmatic Example in Condensed Matter Physics
@@
with $\mathrm{C}\in\left(0,\infty\right]$. The simple choice $\mathrm{C}=\infty$, i.e., $\gamma_{k,q}=\gamma>0$ in (185), is still physically very interesting since, even when $\varepsilon_{k}=0$, the BCS Hamiltonian qualitatively displays most of basic properties of real conventional type I superconductors. See, e.g. [53, Chapter VII, Section 4] or [54] for a more recent mathematically rigorous study.
@@
for $\0c\in\mathbb{C}\backslash\{0\}$, where the ground state energy $E$ is given by
@@
for each $k\in\Lambda_{L}^{\ast}$ and $\0c\in\mathbb{C}\backslash\{0\}$. To get (187), it suffices to inverse the system (189) and substitute the corresponding equalities into (186). Such kind of algebraic manipulation was originally invented for bosons by Bogoliubov [1] in 1947 and it is a key ingredient in the Bogoliubov theory of superfluidity for Helium 4. For fermions, such quadratic Hamiltonians are at the origin of the BCS theory of superconductivity done in 1957 [2, 3, 4] and their diagonalization has been performed in 1958 [5, 59] via a fermionic version of Bogoliubov’s transformation. In modern words, the above algebraic relations are nothing else than the implementation of a Bogoliubov transformation, as given by Theorem 2.8. Compare for instance (189) with Theorem 2.8.
In our framework, it refers to applying Theorem 2.4 and Equation (16) for $E_{0}=\gamma|\Lambda_{L}||c|^{2}$ and operators $\Upsilon_{0}$,$D_{0}$ acting on the $4$-dimensional Hilbert space with orthonormal basis
@@
Avoiding the trivial case $c=0$ for which $D_{0}=0$, such $\mu$ always exists for $\varepsilon_{k}\in\mathbb{R}_{0}^{+}$, $\kappa\in\mathbb{R}$, $\gamma\in\mathbb{R}^{+}$ and $\0c\in\mathbb{C}\backslash\{0\}$. As a consequence, one can satisfy all the assumptions of Theorem 2.4 in this case and use Equations (16) and (17) to deduce in each fiber $\left\{k,-k\right\}$ that
@@
$-8\int_{0}^{\infty}\left\|D_{\tau}\right\|_{2}^{2}\mathrm{d}\tau=\frac{1}{2}\mathrm{tr}\left(\Upsilon_{0}-\sqrt{\Upsilon_{0}^{2}+4D_{0}D_{0}^{\ast}}\right)=2\left(\varepsilon_{k}-\kappa-\sqrt{\left(\varepsilon_{k}-\kappa\right)^{2}+\gamma^{2}|c|^{2}}\right)$
for any $\varepsilon_{k}\in\mathbb{R}_{0}^{+}$, $\kappa\in\mathbb{R}$, $\gamma\in\mathbb{R}^{+}$ and $\0c\in\mathbb{C}\backslash\{0\}$. To recover the results (187)–(188) it then suffices to sum up over the set $\left\{k,-k\right\}_{k\in\Lambda_{L}^{\ast}}$, keeping in mind the explicit form of the Hamiltonian H in Theorem 2.4.
@@
Appendix A Appendix: The Elliptic Operator-Valued Flow
Theorem A.1 (Well-posedness of the flow – Hilbert-Schmidt topology)
The convergence of the flow is a difficult task also solved in [24] under sufficient, albeit still general, conditions. This is done from geometrical characteristics of the flow. We refer here to its ellipticity properties which can first be seen via its constants of motion: For any self-adjoint operator $\Upsilon_{0}$ that is bounded and every non-zero Hilbert-Schmidt operator $D_{0}=\pm D_{0}^{\top}\in\mathcal{L}^{2}(\mathfrak{h})$, one checks that
$\mathrm{tr}\left(\Upsilon_{t}^{2}+4D_{t}D_{t}^{\ast}-\Upsilon_{s}^{2}-4D_{s}D_{s}^{\ast}\right)=0\ ,\qquad s,t\in\mathbb{R}_{0}^{+}\ ,$ (192)
where $\Upsilon_{t}\doteq\Upsilon_{0}+\Delta_{t}$ for any $t\in\mathbb{R}_{0}^{+}$, while $\mathrm{tr}(\cdot)$ denotes the usual trace for operators acting the separable Hilbert space $\mathfrak{h}$. See [24, Equation (23)]. Moreover, if $\Upsilon_{0}D_{0}=D_{0}\Upsilon_{0}^{\top}$ on the domain $\mathcal{D}(\Upsilon_{0}^{\top})\subseteq\mathfrak{h}$ of $\Upsilon_{0}^{\top}$ together with $\Upsilon_{0}\geq-\left(\mu-\varepsilon\right)\mathbf{1}$ for some $\mu\in\mathbb{R}$ and $\varepsilon\in\mathbb{R}^{+}$ and
for some strictly positive constant $\mathrm{C}\in\mathbb{R}^{+}$, then one verifies that, for any $t,s\in\mathbb{R}_{0}^{+}$,
$\Upsilon_{t}D_{t}=D_{t}\Upsilon_{t}^{\top}\qquad\text{and}\qquad\Upsilon_{t}^{2}+4D_{t}D_{t}^{\ast}=\Upsilon_{s}^{2}+4D_{s}D_{s}^{\ast}$ (194)
on the domains $\mathcal{D}(\Upsilon_{0}^{\top})=\mathcal{D}(\Upsilon_{t}^{\top})$ and $\mathcal{D}(\Upsilon_{0}^{2})$, respectively. See [24, Equation (25)]. The last constant of motion can be viewed as the equation of an ellipse in the “commutative” situation.
In the general cases, the natural elipticity properties of the operator-valued flow are written in terms of the operators
$\mathfrak{D}_{t}\doteq\Upsilon_{t}-\mu\mathbf{1}+4D_{t}\left(\Upsilon_{t}^{\top}+\mu\mathbf{1}\right)^{-1}D_{t}^{\ast}\ ,\qquad t\in\mathbb{R}_{0}^{+}\ ,$ (195)
with $\mu\in\mathbb{R}$ such that [^30] Note that $\Upsilon_{0}\geq-\left(\mu-\varepsilon\right)\mathbf{1}$ iff $\Upsilon_{0}^{\top}\geq-\left(\mu-\varepsilon\right)\mathbf{1}$. $\Upsilon_{0}\geq-\left(\mu-\varepsilon\right)\mathbf{1}$ for some $\varepsilon\in\mathbb{R}^{+}$. For instance, in [24, Theorem 3] one proves that the real-valued function
$\zeta\left(t\right)=\inf\sigma\left(\mathfrak{D}_{t}\right)\in\left[-2\mu,\infty\right)\ ,\qquad t\in\mathbb{R}_{0}^{+}\ ,$
$\sigma(\mathfrak{D}_{t})$ being the spectrum of $\mathfrak{D}_{t}$, is upper semicontinuous [^31] For bounded operators $\Upsilon_{0}\in\mathcal{B}\left(\mathfrak{h}\right)$, $\zeta$ is always continuous on $\mathbb{R}_{0}^{+}$. at $t=0$ and continuous on $\mathbb{R}^{+}$ as well as decreasing with $\zeta\left(\mathbb{R}_{0}^{+}\right)\subseteq[0,\zeta\left(0\right)]$ whenever $\zeta\left(0\right)\in\mathbb{R}_{0}^{+}$, while it is continuously increasing with $\zeta\left(\mathbb{R}_{0}^{+}\right)\subseteq[\zeta\left(0\right),0)$ whenever $\zeta\left(0\right)\in\mathbb{R}^{-}$.
Properties of the operator family $(\mathfrak{D}_{t})_{t\in\mathbb{R}_{0}^{+}}$ proven in [24] are important to obtain a good convergence of the operator-valued flow, with possibly exponential decay. This refers to [24, Theorem 4], which corresponds to the following theorem in the Hilbert-Schmidt case:
Theorem A.2 (Asymptotics of the operator-valued flow)
for some $\0\mu\in\mathbb{R}\backslash\{0\}$ and $\varepsilon\in\mathbb{R}^{+}$. Then, as $t\rightarrow\infty$, $D$ exponentially converges in the Hilbert-Schmidt topology (i.e., in $\mathcal{L}^{2}(\mathfrak{h})$) to zero, while $\Delta$ exponentially converges in the trace norm topology (i.e., in $\mathcal{L}^{1}(\mathfrak{h})$) to an operator $\Delta_{\infty}\in\mathcal{L}^{1}(\mathfrak{h})$ satisfying
Compare the assumptions of this theorem with the ones of Theorem 2.4.
Acknowledgments: This work is supported by the Basque Government through the grant IT1615-22 and the BERC 2022-2025 program, by the grant PID2020-112948GB-I00 funded by MCIN/AEI/10.13039/501100011033 and by “ERDF A way of making Europe”. We thank Walter de Siquiera Pedra and Sascha Lill for valuable comments.

### v6

@@ 4 Technical Results: BLS’s Approach to Quadratic Operators
This series is of course well-defined for any $\varphi\in\mathcal{H}$ since $\tilde{V}_{t,s}\varphi$ becomes an absolutely summable series in the Hilbert space $\mathcal{H}$. In fact, $(\tilde{V}_{t,s})_{s,t\in\mathbb{R}}$ is another strongly continuous two-parameter operator family solving the non-autonomous evolution equations (130) and is therefore nothing else that $(V_{t,s})_{s,t\in\mathbb{R}}$, by uniqueness [^29] Compute that, for any $s,t\in\mathbb{R}$ and $\varphi\in\mathcal{H}$, $(\tilde{V}_{t,s}-V_{t,s})\varphi=\int_{s}^{t}\partial_{\tau}\{V_{t,\tau}\tilde{V}_{\tau,s}\}\varphi\ \mathrm{d}\tau=0$. of the solution to (130). This is completely standard. Using the representation (131) of the operator family $(d_{t})_{t\in\mathbb{R}}$ and
@@
Take any self-adjoint operator $h$ satisfying Assumption 1 and use Lemmata 4.1 and 4.2 (i) to represent the group $(\mathrm{e}^{ith})_{t\in\mathbb{R}}$ of Bogoliubov transformations. If $D_{0}\in\mathcal{L}^{2}(\mathfrak{h})$ then $\mathbf{v}_{t}\in\mathcal{L}^{2}(\mathfrak{h})$ for all $t\in\mathbb{R}$.
@@
Take any self-adjoint operator $h$ acting on $\mathcal{H}=\mathfrak{h}\oplus\mathfrak{h}$ satisfying Assumption 1 and use Lemmata 4.1 and 4.2 (i) to represent the group $(\mathrm{e}^{ith})_{t\in\mathbb{R}}$ of Bogoliubov transformations. If $D_{0}\in\mathcal{B}(\mathfrak{h})$ then, for any $t\in\mathbb{R}$,
@@
Proof. Using Lemmata 4.3 and 4.4 with $D_{0}\in\mathcal{B}(\mathfrak{h})$ and $c=\infty$ as well as the mapping (146), we obtain that
@@
Proof. This is an obvious consequence of Lemma 4.9. For more details, see [45].
@@
in the Hilbert-Schmidt topology, as $t\rightarrow 0$. For instance, $\mathcal{O}(1)=0$ when $D_{0}\mathrm{e}^{-i\tau\Upsilon_{0}^{\top}/2}=\mathrm{e}^{i\tau\Upsilon_{0}/2}D_{0}$ for any $\tau\in[0,t]$. In the following lemma, we give the simple case for which the operator $D_{0}$ intertwines [^30] We say that an operator $A$ intertwines between two operators $B$ and $C$ when $AB=CA$. between $\Upsilon_{0}$ and $\Upsilon_{0}^{\top}$ (resp. $\Upsilon_{0}^{\top}$).
@@
Take any self-adjoint operator $h$ acting on $\mathcal{H}=\mathfrak{h}\oplus\mathfrak{h}$ satisfying Assumption 1 with $D_{0}\in\mathcal{B}(\mathfrak{h})$ and use Lemmata 4.1 and 4.2 (i) to represent the group $(\mathrm{e}^{ith})_{t\in\mathbb{R}}$ of Bogoliubov transformations.
@@
    Assume $D_{0}\Upsilon_{0}=-\Upsilon_{0}^{\top}D_{0}$ and $\mathbf{v}_{t}\in\mathcal{L}^{2}(\mathfrak{h})$ for positive times $t$ in a neighborhood of 0. Then, $D_{0}\in\mathcal{L}^{2}(\mathfrak{h})$ and $\mathbf{v}_{t}\in\mathcal{L}^{2}(\mathfrak{h})$ for all $t\in\mathbb{R}$.
@@
    Assume $D_{0}\Upsilon_{0}=\Upsilon_{0}^{\top}D_{0}$ and $\mathbf{v}_{t}\in\mathcal{L}^{2}(\mathfrak{h})$ for positive times $t$ in a neighborhood of 0. If additionally
@@
Step 1: Assume all conditions of the lemma as well as $\Upsilon_{0}D_{0}=\pm D_{0}\Upsilon_{0}^{\top}$. We show here that $D_{0}$ also intertwines between the groups generated by $i\Upsilon_{0}$ and $\pm i\Upsilon_{0}^{\top}$. Since $D_{0}\in\mathcal{B}(\mathfrak{h})$, the equality $\Upsilon_{0}D_{0}=\pm D_{0}\Upsilon_{0}^{\top}$ means in particular that
@@
for small times $t\in\left[0,\delta\right]$. In particular, if $D_{0}\Upsilon_{0}=-\Upsilon_{0}^{\top}D_{0}$ and $\mathbf{v}_{t}\in\mathcal{L}^{2}(\mathfrak{h})$ for positive times $t$ in a neighborhood of 0, then $D_{0}\in\mathcal{L}^{2}(\mathfrak{h})$, which, combined with Lemma 4.8, leads to Assertion (i).
@@
In particular, if (151) holds true as well as $D_{0}\Upsilon_{0}=\Upsilon_{0}^{\top}D_{0}$ and $\mathbf{v}_{t}\in\mathcal{L}^{2}(\mathfrak{h})$ for positive times $t$ in a neighborhood of 0, then $D_{0}\in\mathcal{L}^{2}(\mathfrak{h})$, which, combined with Lemma 4.8, leads to Assertion (ii).
@@
$\Upsilon_{0}\in\mathcal{B}(\mathfrak{h})$ implies the norm continuity of the group $(\mathrm{e}^{it\Upsilon_{0}})_{t\in\mathbb{R}}$ of unitary operators, see, e.g., [31, Chapter I, Proposition 3.5 and Theorem 3.7]. Therefore, Lemma 4.12 (ii) clearly holds also true when $\Upsilon_{0}\in\mathcal{B}(\mathfrak{h})$, $D_{0}\Upsilon_{0}=\Upsilon_{0}^{\top}D_{0}$ and $\mathbf{v}_{t}\in\mathcal{L}^{2}(\mathfrak{h})$ for positive times $t$ in a neighborhood of 0.
@@
Take any self-adjoint operator $h$ satisfying Assumption 1 and use Lemmata 4.1 and 4.2 (i) to represent the group $(\mathrm{e}^{ith})_{t\in\mathbb{R}}$ of Bogoliubov transformations. Assume that $(\mathrm{e}^{ith})_{t\in\mathbb{R}}$ can be implemented in the Fock space $\mathcal{F}$ via a strongly continuous group $(\mathrm{e}^{it\mathrm{H}/2})_{t\in\mathbb{R}}$, $\mathrm{H}=\mathrm{H}^{\ast}$ being a self-adjoint operator acting on $\mathcal{F}$, as in Definition 2.10. Then the following assertions hold true:
@@
The Fock space representation $\mathbf{\pi}_{\mathcal{F}}$ is irreducible, see, e.g. [32, Definition 2.3.7] and [27, Proposition 5.2.2]. It follows from Schur’s lemma [53, Proposition 9.20] that
$\mathbf{\pi}_{\mathcal{F}}\left(\mathcal{A}\right)^{\prime}=\mathbb{C}\,\mathbf{1}\ ,$ (166)
@@
($\partial_{t}^{2}f\left(0\right)$ is in fact the quantum fluctuation of the energy observable $\mathrm{H}$ in the state $\left\langle\Psi,\left(\cdot\right)\Psi\right\rangle_{\mathfrak{h}}$, which is positive by the Cauchy-Schwarz inequality.) As $f\left(t\right)>0$ at least for all $t\in\left[-\delta,\delta\right]$, it follows from (174) that, for any $t\in\left[-\delta,\delta\right]$,
@@
Proof. Combine Proposition 4.14 with Lemma 4.6 as well as Corollaries 4.10 and 4.11.
Appendix A Appendix
A.1 The Elliptic Operator-Valued Flow
Theorem A.1 (Well-posedness of the flow – Hilbert-Schmidt topology)
The convergence of the flow is a difficult task also solved in [24] under sufficient, albeit still general, conditions. This refers to [24, Theorem 4], which corresponds to the following theorem in the Hilbert-Schmidt case:
Theorem A.2 (Asymptotics of the operator-valued flow)
[…]
Compare the assumptions of this theorem with the ones of Theorem 2.4.
A.2 Paradigmatic Example in Condensed Matter Physics
@@
with $\mathrm{C}\in\left(0,\infty\right]$. The simple choice $\mathrm{C}=\infty$, i.e., $\gamma_{k,q}=\gamma>0$ in (182), is still physically very interesting since, even when $\varepsilon_{k}=0$, the BCS Hamiltonian qualitatively displays most of basic properties of real conventional type I superconductors. See, e.g. [54, Chapter VII, Section 4] or [55] for a more recent mathematically rigorous study.
@@
[…]
@@
[…]
In our framework, it refers to applying Theorem 2.4 and Equation (16) for $E_{0}=\gamma|\Lambda_{L}||c|^{2}$ and operators $\Upsilon_{0}$$D_{0}$ acting on the $4$-dimensional Hilbert space with orthonormal basis
@@
$\Upsilon_{\infty}=\sqrt{\Upsilon_{0}^{2}+4D_{0}D_{0}^{\ast}}=\sqrt{\left(\varepsilon_{k}-\kappa\right)^{2}+\gamma^{2}|c|^{2}}\ \mathbf{1}$
@@
$8\int_{0}^{\infty}\left\|D_{\tau}\right\|_{2}^{2}\mathrm{d}\tau=\frac{1}{2}\mathrm{tr}\left(\Upsilon_{0}-\sqrt{\Upsilon_{0}^{2}+4D_{0}D_{0}^{\ast}}\right)=2\left(\varepsilon_{k}-\kappa-\sqrt{\left(\varepsilon_{k}-\kappa\right)^{2}+\gamma^{2}|c|^{2}}\right)$
[…]
@@
Acknowledgments: We thank Walter de Siquiera Pedra, Sascha Lill and an anonymous reviewer for their valuable comments. This work is supported by the Basque Government through the grant IT1615-22 and the BERC 2022-2025 program as well as as the following grants:
  • Grant PID2020-112948GB-I00 funded by MCIN/AEI/10.13039/501100011033 and by “ERDF A way of making Europe”.
  • Grant PID2024-156184NB-I00 funded by MICIU/AEI/10.13039/501100011033 and cofunded by the European Union.